# Optimistic Query Routing in Clustering-based Approximate Maximum Inner Product Search

### v1

@@ 2.2 Understanding the routing behavior
What is left to address is the solution to Problem 1, which is the topic of the remainder of this section. We defer a description of a general approach that uses as much or as little information as available about the data distribution to Appendix B due to space constraints. In the next section, we present a more practical approach that is the foundation of the rest of this work.
2.3 Practical solution via concentration inequalities and sketching
Noting that Problem 1 is captured by the concept of concentration of measure, we resort to results from that literature to find acceptable estimates of $\theta_{i}$’s. In particular, we obtain a solution via a straightforward application of the one-sided Chebyshev’s inequality, resulting in the following lemma.
Lemma 1.
$\theta_{i}=\langle q,\mu_{i}\rangle+\sqrt{\frac{1+\delta}{1-\delta}\;q^{\top}\Sigma_{i}q}.$ (5)
Proof.
$\Pr_{s\sim S_{i}}[s-\langle q,\mu_{i}\rangle\leq\epsilon]\geq 1-\frac{q^{\top}\Sigma_{i}q}{q^{\top}\Sigma_{i}q+\epsilon^{2}}=\frac{1+\delta}{2}\implies\epsilon^{2}=\frac{1+\delta}{1-\delta}\;q^{\top}\Sigma_{i}q.$
Approximating the covariance matrix.
While Equation (5) gives us an algorithm to approximate the threshold $\theta_{i}$, storing $\Sigma_{i}$ can be prohibitive in practice. That is because the size of the matrix grows as $d^{2}$ for each partition. Contrast that with the cost for other routers, such as Mean and NormalizedMean, which only store a single $d$-dimensional vector per partition. To remedy this inefficiency, we reduce the cost of representing $\Sigma_{i}$ by storing a small sketch that approximates it.
Since the procedure we describe is independently applied to each partition, we drop the subscript for $\Sigma_{i}$ and describe the procedure for a single partition. We seek a matrix $\tilde{\Sigma}\in\operatorname*{\mathbb{R}}^{d\times d}$ that approximates $\Sigma$. We define the approximation error as follows:
$\text{err}(\Sigma,\tilde{\Sigma})=\sup_{\begin{subarray}{c}\|v\|_{2}=1\end{subarray}}|v^{\top}\Sigma v-v^{\top}\tilde{\Sigma}v|.$
A standard mechanism to approximate large matrices in order to minimize this error is to compute a low-rank approximation of the original matrix. To that end, let $\Sigma=V\Lambda V^{\top}$ be the eigendecomposition of the positive semi-definite matrix $\Sigma$; where the columns of $V$ contain the orthonormal eigenvectors of $\Sigma$ and $\Lambda$, a $d\times d$ diagonal matrix, contains the non-negative eigenvalues of $\Sigma$ in non-increasing order. The Eckhart-Young-Mirsky Theorem shows that
$\text{err}_{t}(\Sigma)=\inf_{\begin{subarray}{c}\tilde{\Sigma}\in\operatorname*{\mathbb{R}}^{d\times d}\ \text{s.t. rank}(\tilde{\Sigma})=t\end{subarray}}\text{err}(\Sigma,\tilde{\Sigma}),$ (6)
is minimized when $\tilde{\Sigma}=[V\sqrt{\Lambda}]_{t}[V\sqrt{\Lambda}]_{t}^{\top}$, where $[\cdot]_{t}$ corresponds to the operator that selects the first $t$ columns of the matrix.
While (6) is a well-known fact, in practice, it is often the case that the diagonal of the matrix $\Sigma$ contains important information and preserving it fully leads to better approximations of the matrix. As such, we decompose $\Sigma$ as a sum of two matrices: $D$ a diagonal matrix containing the diagonal entries of $\Sigma$, and the residual $R=\Sigma-D$. We preserve the diagonal $D$ fully and approximate $R$ by computing a low-rank approximation of a symmetrization of $R$. Specifically, we decompose
$\Sigma=D+R=D^{\frac{1}{2}}(I+D^{-\frac{1}{2}}RD^{-\frac{1}{2}})D^{\frac{1}{2}},$
and compute a low-rank approximation to $M=D^{-\frac{1}{2}}RD^{-\frac{1}{2}}$. Notice that since $M$ is a symmetric real-valued matrix, it has an eigendecomposition of the form $M=Q\Lambda Q^{\top}$ where $Q$ is a $d\times d$ matrix with orthonormal columns (corresponding to the eigenvectors of $M$) and $\Lambda$ is a $d\times d$ diagonal matrix containing the (possibly negative) eigenvalues of $M$ along the diagonal. Given some target sketch size $t\leq d$, we sketch $\Sigma$ as:
$\tilde{\Sigma}=D+D^{\frac{1}{2}}[Q]_{t}[\Lambda]_{t}[Q]_{t}^{\top}D^{\frac{1}{2}}.$ (7)
While this is not a standard mechanism to sketch PSD matrices and can in the worst case perform worse than the “optimal” low-rank sketch, we show that under certain practical assumptions, this sketch can have error lower than that of standard low-rank approximation.
Lemma 2.
Let $\Sigma\in\operatorname*{\mathbb{R}}^{d\times d}$ be a PSD matrix with diagonal $D$ and the property that $\min_{i\in[d]}{D_{ii}}/{\max_{j\in[d]}D_{jj}}\geq 1-\epsilon$ for some $\epsilon>0$. For every $11\leq t\leq d-$ 1 such that the $(t+1)$-st eigenvalue of $D^{-1/2}\Sigma D^{-1/2}$ is greater than $2$, the sketch $\tilde{\Sigma}$ defined in (7) has the property that $\text{err}(\Sigma,\tilde{\Sigma})\leq\text{err}_{t}(\Sigma)\cdot\frac{1}{1-\epsilon}$.
We provide a proof for the above result in Appendix C and show that for practical datasets, including the ones we use in this work, the assumptions in the lemma are valid.
2.4 The final algorithm
Algorithm 1 Indexing and scoring a single partition with Optimist
1:Set of $n$ points $\mathcal{X}=x_{1},\dots,x_{n}$ defining the partition and target rank $t$ such that $t\leq d$.
2:procedure BuildRouter($\mathcal{X},t$)
3:     Compute and store $\mu\leftarrow\frac{1}{n}\sum_{i=1}^{n}x_{i}$.
4:     Compute $\Sigma\leftarrow\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)(x_{i}-\mu)^{\top}$. $\triangleright$ This does not need to be stored
5:     Decompose $\Sigma$ into diagonal $D$ and residual $R$triangleright $\Sigma=D+R$
6:     Compute eigendecomposition of $D^{-1/2}RD^{-1/2}=Q\Lambda Q^{\top}$
7:     Sort columns of $Q$, $\Lambda$ in non-increasing order.
8:     Store $D$, $\Lambda_{t}\leftarrow[\Lambda]_{t}$ and $Q_{t}\leftarrow[Q]_{t}$
9:Query $q\in\operatorname*{\mathbb{R}}^{d}$ and confidence parameter $\delta>0$.
10:procedure Score($q,\delta$)
11:     Compute $\tilde{q}\leftarrow q\circ D^{1/2}$triangleright Element-wise product with diagonal of $D^{1/2}$
12:     return $\langle q,\mu\rangle+\sqrt{\frac{1+\delta}{1-\delta}\cdot(\|\tilde{q}\|_{2}^{2}+\tilde{q}^{\top}Q_{t}\Lambda_{t}Q_{t}^{\top}\tilde{q})}$
3 Experimental evaluation
In this section, we put our arguments to the test and experimentally evaluate Optimist.
3.1 Setup
Datasets: We use the following suite of benchmark ANN datasets: Text2Image ($10$ million points, $d=200$); Music ($1$m, $d=100$); DeepImage ($10$m, $d=96$); GloVe (1.21.21{.}21.2m, $d=200$); MsMarco-MiniLM (8.88.88{.}88.8m, $d=384$); and, Nq-Ada2 ($2.7$m, $d=1{,}536$). We defer a complete description of these datasets to Appendix A due to space constraints.
Clustering: For our main results, we partition datasets with Spherical KMeans (Dhillon and Modha, 2001). We include in the appendix results from similar experiments but where the clustering algorithm is standard KMeans and Gaussian Mixture Model (GMM). We cluster each dataset into $\sqrt{m}$ shards, where $m$ is the number of data points in the dataset.
Evaluation: Once a dataset is partitioned, we evaluate a router $\tau$ as follows. For each test query, we identify the set of shards to probe using $\tau$. We then perform exact search over the selected shards, obtain the top-$k$ points, and compute recall with respect to the exact top-$k$ set. Because the only source of error is the inaccuracy of the router, the measured recall gauges the effectiveness of $\tau$.
Refer to caption Refer to caption Refer to caption
Figure 2: Top-$k$ recall on Nq-Ada2 versus the number of probed data points. The dataset is partitioned with Spherical KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
Routers: We evaluate the following routers in our experiments:
  • Mean and NormalizedMean: Defined in Equations (2) and (3), respectively;
  • SubPartition $(t)$: Recall that Optimist stores $t+2$ vectors per partition, where $t$ is the rank in Equation (7). We ask if simply partitioning each shard independently into $t+2$ sub-partitions, and recording the sub-partitions’ centroids as the representatives of that shard attains the same routing accuracy as Optimist. At query time, we take the maximum inner product of the query with a shard’s representatives as its score, and sort shards according to this score; and,
  • Optimist $(t,\delta)$: The Optimist router given in Algorithm 1. The parameter $t$ determines the rank of the sketch of the covariance matrix, and $\delta$ determines the degree of optimism. We set $t$ to a maximum of $2\%$ of $d$ but study its effect in Appendix E. $t=d$ indicates that the full covariance matrix is used (i.e., without sketching).
Code: We have implemented all baseline and proposed routers in the Rust programming language. We intend to open-source our code along with experimental configuration to facilitate reproducibility.
3.2 Main results
(a) GloVe
(b) MsMarco-MiniLM
(c) Music
Figure 3: Top-$100$ recall versus the number of probed data points. Setup is the same as Figure 2.
Figure 2 plots top-$k$ recall for $k\in\{1,10,100\}$ versus the percentage of data points examined for Nq-Ada2—Appendix D.1 gives full results for all datasets. Note that, partitioning is by Spherical KMeans, with similar plots for standard KMeans in Appendix D.2 and GMM in Appendix D.3.
We summarize a few key observations. First, among baselines, NormalizedMean generally outperforms Mean and Scann, save for Music where Mean reaches a higher recall. Second, with very few exceptions, Optimist with the full covariance (i.e., $t=d$) does at least as well as NormalizedMean, and often outperforms it significantly. Interestingly, the gap between baselines and Optimist widens as retrieval depth ($k$) increases; a phenomenon that is not surprising.
Finally, while Optimist with $t\ll d$ shows some degradation with respect to the $t=d$ configuration—as anticipated—it still achieves a higher recall than baselines for larger $k$. When $k$ is smaller, the SubPartition router with $t+2$ vectors becomes a strong competitor.
We highlight that, Optimist shines when data points are not on the surface of a sphere (or concentrated close to it). This phenomenon is illustrated in Figure 3, showing top-$100$ recall for a subset of datasets partitioned with Spherical KMeans—refer to Appendix D for results on other datasets.
In particular, on Music, at $95\%$ top-$100$ recall, Optimist with $t\ll d$ needs to probe $54\%$ fewer data points than NormalizedMean; on average Optimist probes $66{,}6$ 666 data points to reach $95\%$ top-$100$ recall whereas NormalizedMean examines $14{,}463$ points. We present the relative savings on all datasets in Table 1. Note that, on DeepImage, no method outperforms NormalizedMean.
[…]
Recall Nq-Ada2 GloVe MsMarco-MiniLM DeepImage Music Text2Image
$90\%$ $18\%$ $11\%$ $22\%$ 00 $38\%$ $23\%$
$95\%$ $20\%$ $5.5\%$ $7.7\%$ 00 $54\%$ $22\%$
We conclude this section by noting that, in Appendix E, we study the effect of $t$ and $\delta$ on the performance of Optimist. We exclude the full discussion from the main prose due to space constraints, but mention that the observations are unsurprising.
3.3 Maximum inner product prediction
We claimed that Optimist is statistically principled. That entails that Optimist should give a more accurate estimate of what the maximum inner product can be for any given query-partition pair. We examine that claim in this section and quantify the prediction error for all routers.
Fix a dataset, whose $C$ partitions are denoted by $\mathcal{P}_{i}$ for $i\in[C]$, along with a router $\tau$ and query $q$. We write $\tau_{i}$ to denote the score computed by $\tau$ for partition $\mathcal{P}_{i}$ and $q$ (e.g., NormalizedMean computes $\langle q,\mu_{i}/\lVert\mu_{i}\rVert_{2}\rangle$ and Optimist gives $\theta_{i}$ per Problem 1).
$\mathcal{E}_{\ell}(\tau,q)=\frac{1}{\ell}\sum_{i=1}^{\ell}\left|\frac{\tau_{\pi_{i}}}{\max_{u\in\mathcal{P}_{\pi_{i}}}\langle q,u\rangle}-1\right|.$ (8)
We measure Equation (8) on all datasets partitioned by Spherical KMeans, and all routers considered in this work. The results are shown in Figure 4, with the remaining datasets in Appendix F, where for each choice of $\ell$, we plot $\mathbb{E}_{q}[\mathcal{E}_{\ell}(\tau,q)]$ using the test query distribution.
[…]
(a) GloVe
(b) MsMarco-MiniLM
(c) Music
Figure 4: Mean prediction error of Equation (8) vs. $\ell$ (percent of total number of shards, $\log$-scale).
4 Concluding remarks
Motivated by our observation that NormalizedMean is an over-estimator of maximum inner product, we formalized the notion of optimism for query routing in clustering-based MIPS and presented a principled optimistic algorithm that estimates the maximum inner product with much greater accuracy and exhibits a more reliable behavior. Results on a suite of benchmark datasets confirm our claims.
We highlight that our algorithm is more suitable for settings where individual shards rest on some external, high-latency storage, so that spending more compute on routing can be tolerated in exchange for fewer amount of data transferred between storage and main memory.
We leave to future work an exploration of a more compact sketching of the covariance matrix; and, an efficient realization of our general solution outlined in Appendix B.
Appendix A Description of datasets
We use the following benchmark datasets in this work:
  • Text2Image: A cross-modal dataset, where data and query points may have different distributions in a shared space [Simhadri et al., 2022]. We use a subset consisting of $10$ million $200$-dimensional data points along with a subset of $10{,}000$ test queries;
  • Music: $1$ million $100$-dimensional points [Morozov and Babenko, 2018] with $1{,}000$ queries;
  • DeepImage: A subset of $10$ million $96$-dimensional points from the billion deep image features dataset [Yandex and Lempitsky, 2016] with $10{,}000$ test queries;
  • GloVe: 1.21.21{.}21.2 million, $200$-dimensional word embeddings trained on tweets [Pennington et al., 2014] with $10{,}000$ test queries;
  • MsMarco-MiniLM: Ms Marco Passage Retrieval v1 [Nguyen et al., 2016] is a question-answering dataset consisting of 8.88.88{.}88.8 million short passages in English. We use the “dev” set of queries for retrieval, made up of $6{,}980$ questions. We embed individual passages and queries using the all-MiniLM-L6-v2 model [^3] Checkpoint at https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. to form a $384$-dimensional vector collection; and,
  • Nq-Ada2: 2.72.72{.}72.7 million, $1{,}536$-dimensional embeddings of the Nq natural questions dataset [Kwiatkowski et al., 2019] with the Ada-002 model. [^4] https://openai.com/index/new-and-improved-embedding-model/
Appendix B General solution
Let $\mathcal{D}$ be an unknown distribution over $\operatorname*{\mathbb{R}}^{d}$ from which shard $\mathcal{P}_{i}$ is sampled. It is clear that, if we are able to accurately approximate the quantiles of $S_{i}$, we can obtain an estimate $\theta_{i}$ satisfying Problem 1.
@@ Appendix B General solution
$\theta_{i}=\langle q,\mu_{i}\rangle+\sqrt{q^{\top}\Sigma_{i}q}\cdot\Phi_{\mathcal{N}(0,1)}^{-1}\left(\frac{1+\delta}{2}\right),$ (9)
@@
Notice in the above example that, we first modeled the moments of $S_{i}$ using the moments of $\mathcal{D}$. We then approximated the CDF (or equivalently, the quantile function) of the distribution of $S_{i}$ using its moments—in the Gaussian case, the approximation with the first two moments is, in fact, exact.
@@
In our second step, we wish to find a distribution $\tilde{S}_{i}$ such that $m_{j}(\tilde{S}_{i})\approx m_{j}(S_{i})$ for all $j\in[r]$. This can be done using the “method of moments,” a classic technique in statistics Pearson [1936] that offers guarantees [Kong and Valiant, 2017, Braverman et al., 2022] in terms of a distributional distance such as the Wasserstein-1 distance.
@@
While using higher-order moments can lead to a better approximation, computing the $j$-th moment for $j>2$ can be highly expensive considering the dimensionality of datasets seen in practice, as the memory requirement to store the tensor $m_{j}(\mathcal{D})$ grows as $d^{j}$. Hence, we leave exploration and design of an efficient version of this two-step approach as future work.
Appendix C Sketching the covariance matrix

### v3

@@ 2.2. Understanding the routing behavior
What is left to address is the solution to Problem 1, which is the topic of the remainder of this section. We first describe a general approach that uses as much or as little information as available about the data distribution. We then present a more practical approach that is the foundation of the rest of this work.
2.3. General solution
Let $\mathcal{D}$ be an unknown distribution over $\operatorname*{\mathbb{R}}^{d}$ from which $\mathcal{P}_{i}$ is sampled. It is clear that, if we are able to accurately approximate the quantiles of $S_{i}$, we can obtain an estimate $\theta_{i}$ satisfying Problem 1.
@@ 2.3. General solution
(5) $\theta_{i}=\langle q,\mu_{i}\rangle+\sqrt{q^{\top}\Sigma_{i}q}\cdot\Phi_{\mathcal{N}(0,1)}^{-1}\left(\frac{1+\delta}{2}\right),$
@@
Notice in the above example that, we first modeled the moments of $S_{i}$ using the moments of $\mathcal{D}$. We then approximated the CDF (equivalently, the quantile function) of the distribution of $S_{i}$ using its moments—in the Gaussian case, the approximation with the first two moments is, in fact, exact.
@@
In our second step, we wish to find a distribution $\tilde{S}_{i}$ such that $m_{j}(\tilde{S}_{i})\approx m_{j}(S_{i})$ for all $j\in[r]$. This can be done using the “method of moments,” a classic technique in statistics (Pearson, 1936) that offers guarantees (Kong and Valiant, 2017; Braverman et al., 2022) in terms of a distributional distance such as the Wasserstein-1 distance.
@@
While using higher-order moments can lead to a better approximation, computing the $j$-th moment for $j>2$ can be highly expensive considering the dimensionality of datasets seen in practice, as the memory requirement to store the tensor $m_{j}(\mathcal{D})$ grows as $d^{j}$. We leave exploration and design of an efficient version of this two-step approach as future work.
2.4. Practical solution via concentration inequalities and sketching
Noting that Problem 1 is captured by the concept of concentration of measure, we resort to results from that literature to find acceptable estimates of $\theta_{i}$’s. In particular, we obtain a solution via an application of the one-sided Chebyshev’s inequality, resulting in the following lemma.
Lemma 0.
(6) $\theta_{i}=\langle q,\mu_{i}\rangle+\sqrt{\frac{1+\delta}{1-\delta}\;q^{\top}\Sigma_{i}q}.$
Proof.
$\Pr_{s\sim S_{i}}[s-\langle q,\mu_{i}\rangle\leq\epsilon]$ $\geq 1-\frac{q^{\top}\Sigma_{i}q}{q^{\top}\Sigma_{i}q+\epsilon^{2}}=\frac{1+\delta}{2}$
$\implies\epsilon^{2}=\frac{1+\delta}{1-\delta}\;q^{\top}\Sigma_{i}q.$
While Equation (6) gives us an algorithm to approximate the threshold $\theta_{i}$, storing $\Sigma_{i}$ can be prohibitive in practice. That is because the size of the matrix grows as $d^{2}$ for each partition. Contrast that with the cost for other routers, such as Mean and NormalizedMean, which only store a single $d$-dimensional vector per partition (i.e., the mean vector). To remedy this inefficiency, we design a compact approximation of $\Sigma_{i}$ that can be plugged into Equation (6) to replace $\Sigma_{i}$. We describe that next to complete our algorithm.
2.4.1. Approximating the covariance matrix.
Since the procedure we describe is independently applied to each partition, we drop the subscript in $\Sigma_{i}$ and focus on a single partition.
We seek a matrix $\Sigma^{\ast}\in\operatorname*{\mathbb{R}}^{d\times d}$ that approximates the positive semi-definite (PSD) matrix $\Sigma$, by minimizing the following standard approximation error:
(7) $\text{err}(\Sigma,\Sigma^{\ast})\triangleq\sup_{\begin{subarray}{c}\|v\|_{2}=1\end{subarray}}|v^{\top}\Sigma v-v^{\top}\Sigma^{\ast}v|.$
As we seek a compact approximation, we constrain the solution space to rank-$t$ matrices: We require $\text{rank}(\Sigma^{\ast})=t$ for some $t\ll d$.
A standard solution to this problem is a result of the Eckhart-Young-Mirsky Theorem: Let $\Sigma=V\Lambda V^{\top}$ be the eigendecomposition of $\Sigma$. Then, $\Sigma^{\ast}=[V\sqrt{\Lambda}]_{t}[V\sqrt{\Lambda}]_{t}^{\top}$, where $[\cdot]_{t}$ is the operator that selects the first $t$ columns of its argument, optimally approximates $\Sigma$ under the stated constraints. We denote this solution by $\Sigma^{\textsc{lr}}_{t}$.
We could stop here and use $\Sigma^{\textsc{lr}}_{t}$ in lieu of $\Sigma$ in Equation (6). However, while that would be the optimal choice in the general case, for real-world datasets, we show that we can find a better approximation. In particular, we present the following:
Definition 2 (Masked Sketch of Rank $t$).
Rewrite $\Sigma$ as follows, with $D$ containing its diagonal and $R=\Sigma-D$:
$\Sigma=D+R=D^{\frac{1}{2}}(I+\underbrace{D^{-\frac{1}{2}}RD^{-\frac{1}{2}}}_{R_{\circ}})D^{\frac{1}{2}}.$
Let $Q\Lambda Q^{\top}$ be the eigendecomposition of $R_{\circ}$. We call the following the Masked Sketch of Rank $t$ of $\Sigma$:
(8) $\Sigma^{\textsc{ms}}_{t}=D+D^{\frac{1}{2}}[Q]_{t}[\Lambda]_{t}[Q]_{t}^{\top}D^{\frac{1}{2}}.$
The following result establishes that, under certain practical assumptions, $\text{err}(\Sigma,\Sigma^{\textsc{ms}}_{t})$ can be lower than $\text{err}(\Sigma,\Sigma^{\textsc{lr}}_{t})$.
Lemma 0.
Let $\Sigma\in\operatorname*{\mathbb{R}}^{d\times d}$ be a PSD matrix with diagonal $D$ for which $\min_{i\in[d]}{D_{ii}}/{\max_{i\in[d]}D_{ii}}\geq 1-\epsilon$ for some $\epsilon>0$. For every $11\leq t\leq d-$ 1 such that the $(t+1)$-th eigenvalue of $D^{-1/2}\Sigma D^{-1/2}$ is greater than $1$, we have that: $\text{err}(\Sigma,\Sigma^{\textsc{ms}}_{t})\leq\text{err}(\Sigma,\Sigma^{\textsc{lr}}_{t})\cdot\frac{1}{1-\epsilon}$.
We provide a proof for the above result in the next section and show that for real-world datasets used in this work, the assumptions in the lemma hold. But before we do so, we wish to recap our final solution.
Algorithm 1 Indexing and scoring a single partition with Optimist
1:Partition $\mathcal{P}$ and target rank $t\ll d$.
2:procedure BuildRouterForPartition($\mathcal{P},t$)
3:     $\mu\leftarrow\frac{1}{\lvert\mathcal{P}\rvert}\sum_{u\in\mathcal{P}}u$
4:     $\Sigma\leftarrow\frac{1}{\lvert\mathcal{P}\rvert}\sum_{u\in\mathcal{P}}(u-\mu)(u-\mu)^{\top}$
5:     $R_{\circ}=D^{-1/2}(\Sigma-D)D^{-1/2}$triangleright $D$ is the diagonal of $\Sigma$
6:     Find eigendecomposition $R_{\circ}=Q\Lambda Q^{\top}$
7:     Sort columns of $Q$, $\Lambda$ in non-increasing order
8:     $\Lambda_{t}\leftarrow[\Lambda]_{t}$, $Q_{t}\leftarrow[Q]_{t}$
9:     return $\{\mu,D,\Lambda_{t},Q_{t}\}$
10:Query $q\in\operatorname*{\mathbb{R}}^{d}$ and optimism parameter $\delta>0$.
11:procedure ScorePartitionForQuery($q,\delta$; $\{\mu,D,\Lambda_{t},Q_{t}\}$)
12:     $\tilde{q}\leftarrow q\circ\text{diag}(D^{1/2})$triangleright Element-wise product
13:     return $\langle q,\mu\rangle+\sqrt{\frac{1+\delta}{1-\delta}\cdot(\|\tilde{q}\|_{2}^{2}+\tilde{q}^{\top}Q_{t}\Lambda_{t}Q_{t}^{\top}\tilde{q})}$
2.4.2. The final algorithm
2.5. Proofs and justification of assumptions